# Decay-protected superconducting qubit with fast control enabled by integrated on-chip filters

### v1

@@ Abstract
Subharmonic control of superconducting qubits simultaneously facilitates fast single-qubit gates, long coherence times, and reduced thermal load. However, its implementation requires a very strong coupling of the drive line to the qubit or an excessively strong microwave signal. To address this challenge, we introduce on-chip filters of the qubit drive that exhibit a stopband at the resonance frequency and ultra-strong coupling at the subharmonic frequency. The on-chip filters are designed to demonstrate qubit $T_{1}$ in the millisecond range while enabling fast gates of tens of nanoseconds with subharmonic control. For a resonant drive, we experimentally show a maximum of 200-fold improvement in the measured $T_{1}$ and a factor of 220 suppression in the measured Rabi frequency at the stopband. Furthermore, we successfully implemented subharmonic driving of Rabi oscillations with a $\pi$-pulse duration of 12 ns. Our demonstration of on-chip filters and the resulting introduction of efficient subharmonic driving in a two-dimensional quantum processor paves the way for a scalable qubit architecture with reduced thermal load and noise from the control line.
@@ Introduction
Superconducting qubits, in particular, have exhibited remarkable progress, with relaxation times extending from a few nanoseconds [14, 18] to nearly a few milliseconds [19, 20, 21, 22] over the past two decades. The quest for further improvement of the coherence time is still ongoing. Alternatively, gates can be sped up to further improve the CT2GT ratio [23, 24]. However, this requires strong external coupling of the qubit to the drive line, which inadvertently decreases the coherence of the qubit, thereby posing a significant trade-off.
The coherence times of a qubit are influenced not only by the external coupling but also by intrinsic losses which in properly designed qubits, represent the primary loss mechanism, thus limiting the qubit performance. These losses arise from various sources, such as the presence of two-level-system defects near the qubit [25, 26, 27, 28, 29, 30, 31, 32] associated with dielectric losses and quasiparticle tunneling [33, 34, 35, 36]. Remarkable progress has been made in understanding [37, 38, 39, 40, 41, 42, 43, 44, 45] and mitigating these internal losses through the implementation of high-coherence materials [46, 47, 48, 49, 21, 22], fabrication recipes [50, 19, 20, 51], and improved qubit geometry [52, 53], leading to a situation where the external coupling should be made very weak not to limit qubit coherence, thus calling for high drive power at the qubit.
@@
In this work, we introduce designs and provide implementations of on-chip filters for subharmonic driving in two-dimensional qubit schemes. The on-chip filters are devised to fully isolate the qubit from the drive line at its resonance frequency while establishing two orders of magnitude stronger coupling at the subharmonic frequency compared to standard drive lines. This result enables fast single-qubit gates via subharmonic drive while demonstrating the $T_{1}$ limit owing to the external coupling of milliseconds, thereby substantially improving the corresponding CT2GT ratio. Encouragingly, the introduced on-chip filters facilitate the use of over 60 dB of attenuation in the control lines to ensure a thermal noise photon below $2.5\times 10^{-3}$. Furthermore, we show that the use of a bulky low-pass filter creates unwanted resonances in the qubit drive line, modulating the measured $T_{1}$ in extreme ranges. Thus our first demonstration of subharmonic driving of a qubit in two dimensions, integrated with an on-chip filter, seems a promising pathway for scalable superconducting quantum processing units.
@@ Results
We have developed two distinct versions of on-chip filters: $\mathrm{\lambda/4}$ and $\mathrm{\lambda/2}$ filters as shown in Fig. 1a, b. Essentially, these filters are coplanar-waveguide (CPW) transmission lines that connect capacitively to the qubit at the position $x=0$ and end in an open circuit at $x=L_{\mathrm{f}}$, where $L_{\mathrm{f}}$ is the length of the filter. In the case of the $\mathrm{\lambda/4}$ filter, $L_{\mathrm{f}}=\lambda_{\mathrm{q}}/4$, where $\lambda_{\mathrm{q}}$ is the wavelength of the mode in the drive line at the qubit frequency. Similarly, for the $\mathrm{\lambda/2}$ filter, $L_{\mathrm{f}}=\lambda_{\mathrm{q}}/2$. We also note that $\mathrm{\lambda/2}$ filter couples to the qubit at two separate positions as opposed to the single-point coupling in the case of the $\mathrm{\lambda/4}$ filter.
@@
We target the length of the on-chip filters such that their stopband is centered around 5 GHz. The results of our numerical simulations of these filters are shown in Fig. 2. For a detailed description of the electromagnetic (EM) simulations, we refer to the Supplementary Information S1.
Figure 2: Simulated transmission and relaxation time of the qubit from drive lines. a Schematic of a 3-port network labeled P1, P2, and P3. Port P1 is connected to a qubit drive line through a filter which is a standard transmission line for Q1, $\lambda/4$ filter for Q2, and $\lambda/2$ filter for Q3. b Equivalent lumped-element model, where $\mathrm{C_{d}}$ is the capacitance between the qubit and the drive line, $\mathrm{C_{g}}$ is the capacitance between the qubit and the readout resonator, and $\mathrm{C_{q}}$ is the shunt capacitance of the qubit to the ground. The coupling capacitances $\mathrm{C_{d}}$ are 80 aF, 4.5 fF, and 9.2 fF for qubits Q1, Q2, and Q3, respectively. c Transmission amplitude from the qubit port to the drive line, $S_{\mathrm{21}}$, as a function of frequency for the different filter configurations as indicated. The $\lambda/4$ and $\lambda/2$ filters are designed to have a stopband at 5 GHz. The sharp features around 7 GHz correspond to the individual readout resonator frequencies for each qubit. The inset highlights the region in the vicinity of the filter stopband for convenient comparison. d Estimated $T_{\mathrm{1}}$ of the qubit owing to its coupling just to the drive line as a function of qubit frequency. This estimation is based on the total capacitance of the qubit and the input admittance $Y_{\mathrm{in}}$ parallel to the qubit port P2. At the stopband, the estimated relaxation times of the qubit for the standard drive, the $\lambda/4$ filter, and for the $\lambda/2$ filter are approximately 0.3 ms, 7161 ms, and 2421 ms, respectively.
@@
Figures 2a, b depict a schematic and its equivalent lumped-element model of a qubit coupled to an on-chip filter and to a readout resonator. In the model, 50-$\Omega$ lumped ports, labeled P1 for the drive line, P2 for the qubit, and P3 for the readout resonator, are introduced for EM simulations. We focus on the scattering parameter $S_{\mathrm{21}}$, which relates to the power transmission from the drive line to the qubit port, and on the input admittance $Y_{\mathrm{in}}$ parallel to the qubit. Here, $Y_{\mathrm{in}}$ is calculated when port P1 is matched to an impedance of 50-$\Omega$, while port P3 is grounded. With this procedure, one can estimate the external coupling rate just from the drive line to be $\gamma_{\mathrm{q}}(\omega)=\mathrm{Re}[Y_{\mathrm{in}}(\omega)]/C_{\mathrm{q}}$ [59, 60]. In the low-temperature limit, the relaxation time can be simply calculated using $T_{\mathrm{1}}(\omega)=1/\gamma_{\mathrm{q}}(\omega)$, where we assume negligible internal losses of the qubit.
@@
Figure 2c shows that within the filter stopband centered at 5 GHz, the transmission from the qubit port to the transmission line lies below $-110$ dB, greatly suppressed from $-72$ dB obtained for the standard drive line. Importantly, within the subharmonic frequency range of 1–2 GHz, we observe a notable increase of over 30 dB in the transmission of the on-chip filters over the standard case. The observed sharp features around 7 GHz correspond to the resonances from the readout resonators.
@@
In Fig. 2d, we examine the relaxation time $T_{\mathrm{1}}$ of the qubit resulting from the external coupling to the drive line. Within the filter stopband, we observe a substantial enhancement in $T_{\mathrm{1}}$ by a factor of over 1000 with the $\lambda/4$ and $\lambda/2$ filters compared to the standard drive line. The estimated $T_{\mathrm{1}}$ times at 5 GHz are as follows: 0.3 ms for the standard drive, 7161 ms for $\lambda/4$ filter, and  2421 ms for $\lambda/2$ filter. The widths of the frequency ranges where qubit $T_{1}$ exceeds 1 ms are 70 MHz and 450 MHz for $\lambda/4$ and $\lambda/2$ filter, respectively. We attribute this difference of feasible frequency ranges to the fact that the $\lambda/2$ filter is of higher order than the $\lambda/4$ filter since the coupling points of the qubit are located at the voltage anti-nodes where the gradient of the voltage with respect to position vanishes.
@@
Assuming a coupling that results in $T_{1}=1$ ms for a qubit driven resonantly through a standard transmission line, we estimate that a microwave pulse with a peak power of $-11$ dBm at room temperature is necessary to implement 10-ns single-qubit gates. With the above simulated parameters of the on-chip filters, corresponding peak powers of $-16$ dBm and $-22$ dBm will be required for $\lambda/4$ filter and $\lambda/2$ filter, respectively, using a subharmonic drive approach. Thus the filters combined with subharmonic driving may yield a significant reduction in power consumption for single-qubit gates. These values are calculated considering 60 dB of attenuation in the drive line and a simple rectangular pulse shape. Note that the peak powers may significantly increase for complex pulse shapes which are necessary to minimize the leakage out of the computational subspace for high-fidelity gates [23].
@@
Figure 3: Characterization of the on-chip filters. a,b Measured (dots) and simulated (dashed line) Rabi frequencies of the qubit for (a) $\lambda/4$ and (b) $\lambda/2$ filters as functions of the qubit frequency $f_{\mathrm{q}}=\omega_{\mathrm{q}}/(2\pi)$. We swept qubit frequency in the range of 4 to 6.5 GHz and carried out the Rabi measurement to characterize the coupling strength of the drive lines to the qubit. The dashed line represents the Rabi frequency estimated using $f_{\mathrm{R}}=\Omega_{\mathrm{R}}/2\pi=2\sqrt{\gamma_{\mathrm{q}}}\beta$, where $\gamma_{\mathrm{q}}=\mathrm{Re}[Y_{\mathrm{in}}]/C_{\mathrm{q}}$ and $\beta\propto\omega_{\mathrm{q}}^{-1/2}V$, where V is the peak-to-peak voltage amplitude at the chip level. Note that the measured Rabi frequency for the $\lambda/2$ filter shows two distinct stopbands centered around 4.7 GHz and 5.7 GHz, the presence of which is attributed to the unequal coupling at positions $x=0$ and $x=\mathrm{\lambda/2}$. c,d Measured (dots) and simulated (dashed line) $T_{1}$ of the qubit for (c) $\lambda/4$ and (d) $\lambda/2$ filters as functions of the qubit frequency. In the simulation, the dielectric loss tangent of $3\times 10^{-6}$ obtained from the fit is used to account for dielectric losses limiting the $T_{1}$. The error bars represent $1\sigma$ fitting uncertainty.
@@
We follow the typical steps to fabricate the device on a silicon wafer, which includes a niobium metallization layer on top. The micron-sized features in the niobium are created using maskless lithography and etching, while the Manhattan-style Josephson junctions are patterned using electron-beam lithography and thermal evaporation of aluminium. We cool down the device in a commercial Bluefors XLD500 dilution refrigerator for cryogenic measurements. For more detailed information on the fabrication process and the experimental setup, see the Supplementary Information S2 & S3.
@@
We carry out typical time-domain measurements to characterize the device parameters summarized in Table 1 of the Supplementary Information S4. To quantify the performance of the on-chip filters, we measured the Rabi frequency and the $T_{1}$ of the qubit around the target filter frequency of 5 GHz. A comprehensive description of the measurement sequence can be found in greater detail in the Supplementary Information S4.
@@
Figures 3a, b show the measured Rabi frequency for qubits driven through the $\mathrm{\lambda/4}$ and $\mathrm{\lambda/2}$ filters. We observe an impressive 50-fold and 220-fold suppression in measured Rabi frequency at the stopbands of $\lambda/4$ and $\lambda/2$ filters, respectively, corresponding to 34 dB and 47 dB changes in $S_{21}$. The simulated data depicted in Fig. 3a, b correspond to the Rabi frequency derived from the input admittance $Y_{\mathrm{in}}$ obtained through EM simulations, see Methods. Conveniently, the measured and simulated Rabi frequencies exhibit a high level of agreement. Some qualitative deviations observed can be attributed to parasitic resonance and imperfections stemming from coupling with the control lines.
@@
Note the significant discrepancy in the measured Rabi frequency for the $\lambda/2$ filter shown in Fig. 3b in comparison to the simulated transmission depicted in Fig. 2c. In contrast to a single stopband at 5 GHz, we observe two stopbands centered around 4.7 GHz and 5.7 GHz. As detailed in Supplementary Information S1, this splitting of the stop band mainly arises from the asymmetry in the coupling capacitance between the drive line and the qubit at the positions $x=0$ and $x=\lambda/2$, in addition to the fact that the coupling region at $x=\lambda/2$ is of finite length. Consequently, we show a good agreement between the measured and simulated results in Fig. 3b by introducing an asymmetric coupling at the open-end of the $\mathrm{\lambda/2}$ filter in the EM simulation.
@@
Figures 3c, d depict the measured $T_{1}$ for the $\mathrm{\lambda/4}$ and $\mathrm{\lambda/2}$ filter. In our device, we consistently observe a $T_{1}$ ceiling of approximately 10 $\mu\mathrm{s}$ across all qubits. To investigate the reason for this limited $T_{1}$, we fit the Purcell decay and dielectric-loss models to the measured $T_{1}$ of qubit Q1 near its readout resonator frequency, see Supplementary Information S5. From the fit, we obtain a dielectric loss tangent of $3\times 10^{-6}$, confirming dielectric loss as the primary loss mechanism in our device. Consequently, we used the dielectric loss tangent of $3\times 10^{-6}$ in the EM simulations, yielding the simulation results for qubits Q2 and Q3 in good agreement with the experiments.
@@
For the qubit $T_{1}$ measured at the sweet spot, we observed a remarkable 20-fold improvement for the $\mathrm{\lambda/4}$ filter and an even more substantial 200-fold improvement for the $\mathrm{\lambda/2}$ filter in comparison to the stadard case of no on-chip filter. This enhancement has the potential for further improvements through the use of low-loss dielectric materials and implemention of state-of-the-art fabrication processes for high coherence.
@@
Figures 4a,b present the flux spectroscopy results for qubit Q3 coupled to $\lambda/2$ filter using both resonant and subharmonic driving techniques. By sweeping the voltage across a resistor installed in the flux line at room temperature, we tune the frequency of qubit Q3 in a range of 500 MHz near the second stopband of the $\lambda/2$ filter centered around 5.7 GHz. With resonant driving, we observe a major reduction in the qubit linewidth as it approaches the stopband at 5.7 GHz. Note that also the two-photon process to excite the qubit from its ground state to the second excited state vanishes at different flux bias but frequency matching to 5.7 GHz, consistent with the filter protecting the qubit at its stopband rather than the qubit itself being substantially different at a certain flux bias. However, with subharmonic driving, we observe a broadened qubit linewidth at around $f=5.7/3=1.9$ GHz. This observation indicates that the $\lambda/2$ filter exhibits strong coupling at the subharmonic frequency but very weak coupling at the resonance frequency, which agrees with our intended design and purpose.
Figure 4: Spectroscopy and Rabi oscillations of qubit Q3 with resonant driving and subharmonic driving through the $\mathrm{\lambda/2}$ filter. a,b Normalized readout signal of the qubit as a function of the flux bias and the frequency of the (a) resonant drive and (b) subharmonic drive. The top diagonal feature indicates the lowest transition frequency of the qubit and the bottom feature reveals the two-photon transition to the second excited state of the qubit. c,d Normalized readout signal of the qubit as a function of the driving-voltage amplitude at room temperature and the frequency of the (c) resonant drive and (d) subharmonic drive for a flux bias of 0.32 V. Note the power broadening and up to $-70$ MHz alternating-current (AC) Stark shift in (d) with increasing pulse amplitude. e,f Normalized readout signal of the qubit as a function of the pulse length and the frequency of the (e) resonant drive and (f) subharmonic drive for a flux bias of 0.32 V. In (e), the pulse amplitude 0.55 V yields a Rabi frequency of 17.57 MHz and in (f), we have 0.9 V and 31.12 MHz, respectively, at the Stark-shifted frequency of 1.809 GHz.
@@
Figures 4d,e illustrate Rabi oscillations of the qubit under resonant and subharmonic driving. For room-temperature pulse amplitudes of 0.55 V and 0.9 V, the measured Rabi frequencies for resonant and subharmonic driving are 17.57 MHz and 31.12 MHz, respectively. In the case of subharmonic drive, the Rabi frequency corresponding to the slowest oscillation occurs with an AC Stark shift of $-57.67$ MHz.

### v2

@@ Abstract
Achieving fast gates and long coherence times for superconducting qubits presents challenges, typically requiring either a stronger coupling of the drive line or an excessively strong microwave signal to the qubit. To address this, we introduce on-chip filters of the qubit drive exhibiting a stopband at the qubit frequency, thus enabling long coherence times and strong coupling at the subharmonic frequency, facilitating fast single-qubit gates, and reduced thermal load. The filters exhibit an extrinsic relaxation time of a few seconds while enabling sub-10-ns gates with subharmonic control. Here we show up to 200-fold improvement in the measured relaxation time at the stopband. Furthermore, we implement subharmonic driving of Rabi oscillations with a $\pi$ pulse duration of 12 ns. Our demonstration of on-chip filters and efficient subharmonic driving in a two-dimensional quantum processor paves the way for a scalable qubit architecture with reduced thermal load and noise from the control line.
@@ Introduction
Superconducting qubits, in particular, have exhibited significant progress, with relaxation times extending from a few nanoseconds [14, 18, 19, 20, 21, 22] to nearly a few milliseconds [23, 24, 25, 26] over the past two decades. The quest for further improvement of the coherence time is still ongoing. Alternatively, gates can be sped up to further improve the CT2GT ratio [27, 28]. However, this requires strong external coupling of the qubit to the drive line, which inadvertently decreases the coherence of the qubit, thereby posing a significant trade-off.
The coherence times of a qubit are influenced not only by the external coupling but also by intrinsic losses which in properly designed qubits, represent the primary loss mechanism, thus limiting the qubit performance. These losses arise from various sources, such as the presence of two-level-system defects near the qubit [29, 30, 31, 32, 33, 34, 35, 36] associated with dielectric losses and quasiparticle tunneling [37, 38, 39, 40]. Significant progress has been made in understanding [41, 42, 43, 44, 45, 46, 47, 48, 49] and mitigating these internal losses through the implementation of high-coherence materials [50, 51, 52, 53, 25, 26], fabrication recipes [54, 23, 24, 55], and improved qubit geometry [56, 57], leading to a situation where the external coupling should be made very weak not to limit qubit coherence, thus calling for high drive power at the qubit.
@@
In this work, we introduce designs and provide implementations of on-chip filters for subharmonic driving in two-dimensional qubit schemes. The on-chip filters are devised to fully isolate the qubit from the drive line at its resonance frequency while establishing two orders of magnitude stronger coupling at the subharmonic frequency compared to standard drive lines. This result enables fast single-qubit gates via subharmonic drive while demonstrating the $T_{1}$ limit owing to the external coupling of seconds, thereby substantially improving the corresponding CT2GT ratio. Encouragingly, the introduced on-chip filters facilitate the use of over 60 dB of attenuation in the control lines to ensure a thermal noise photon below $2.5\times 10^{-3}$. Furthermore, we show that the use of a bulky low-pass filter creates unwanted resonances in the qubit drive line, modulating the measured $T_{1}$ in extreme ranges. Thus our demonstration of subharmonic driving of a qubit in two dimensions, integrated with an on-chip filter, seems a promising pathway for scalable superconducting quantum processing units.
@@ Results
We have developed two distinct versions of on-chip filters: $\mathrm{\lambda/4}$ and $\mathrm{\lambda/2}$ filters as shown in Fig. 1a, b. Essentially, these filters are coplanar-waveguide (CPW) transmission lines that connect capacitively to the qubit at the position $x=0$ and end in an open circuit at $x=L_{\mathrm{f}}$, where $L_{\mathrm{f}}$ is the length of the filter. In the case of the $\mathrm{\lambda/4}$ filter, $L_{\mathrm{f}}=\lambda_{\mathrm{q}}/4$, where $\lambda_{\mathrm{q}}$ is the wavelength of the mode in the drive line at the qubit frequency. Similarly, for the $\mathrm{\lambda/2}$ filter, $L_{\mathrm{f}}=\lambda_{\mathrm{q}}/2$. We also note that $\mathrm{\lambda/2}$ filter couples to the qubit at two separate positions as opposed to the single-point coupling in the case of the $\mathrm{\lambda/4}$ filter. This two-point coupling design is inspired from multi-point coupling investigated in waveguide quantum electrodynamics [61, 62].
Figure 2: Simulated transmission and relaxation time of the qubit from drive lines. a Schematic of a 3-port network labeled P1, P2, and P3. Port P1 is connected to a qubit drive line through a filter which is a standard transmission line for Q1, $\lambda/4$ filter for Q2, and $\lambda/2$ filter for Q3. b Equivalent lumped-element model, where $\mathrm{C_{d}}$ is the capacitance between the qubit and the drive line, $\mathrm{C_{g}}$ is the capacitance between the qubit and the readout resonator, and $\mathrm{C_{q}}$ is the shunt capacitance of the qubit to the ground. The coupling capacitances $\mathrm{C_{d}}$ are 80 aF, 4.5 fF, and 9.2 fF for qubits Q1, Q2, and Q3, respectively. c Transmission amplitude from the qubit port to the drive line, $S_{\mathrm{21}}$, as a function of frequency for the different filter configurations as indicated. The $\lambda/4$ and $\lambda/2$ filters are designed to have a stopband at 5 GHz. The inset highlights the region in the vicinity of the filter stopband for convenient comparison. The shaded gray region corresponds to transmission at subharmonic frequency. d Estimated relaxation time, $T_{\mathrm{1,ext}}$, of the qubit owing to its coupling just to the drive line as a function of qubit frequency. This estimation is based on the total capacitance of the qubit and the input admittance $Y_{\mathrm{in}}$ parallel to the qubit port P2. At the stopband, the estimated relaxation times of the qubit for the standard drive, the $\lambda/4$ filter, and for the $\lambda/2$ filter are approximately 0.3 ms, 7161 ms, and 2421 ms, respectively.
@@
We target the length of the on-chip filters such that their stopband is centered around 5 GHz. The results of our numerical simulations of these filters are shown in Fig. 2. For a detailed description of the electromagnetic (EM) simulations, we refer to the Supplementary Note S1 in Supplementary Information.
@@
Figures 2a, b depict a schematic and its equivalent lumped-element model of a qubit coupled to an on-chip filter and to a readout resonator. In the model, 50-$\Omega$ lumped ports, labeled P1 for the drive line, P2 for the qubit, and P3 for the readout resonator, are introduced for EM simulations. We focus on the scattering parameter $S_{\mathrm{21}}$, which relates to the power transmission from the drive line to the qubit port, and on the input admittance $Y_{\mathrm{in}}$ parallel to the qubit. Here, we are only interested in the coupling between the drive line and the qubit. Therefore, both $S_{21}$ and $Y_{\mathrm{in}}$ are calculated for the drive line port P1 matched to an impedance of 50 $\Omega$ and the resonator port P3 grounded. With this procedure, one can estimate the external coupling rate just from the drive line to be $\gamma_{\mathrm{q}}(\omega)=\mathrm{Re}[Y_{\mathrm{in}}(\omega)]/C_{\mathrm{q}}$ [65, 66]. In the low-temperature limit, the relaxation time is given by $T_{\mathrm{1}}(\omega)=[1/T_{\mathrm{1,ext}}(\omega)+1/T_{\mathrm{1,diel}}(\omega)+1/T_{\mathrm{1,oth}}(\omega)]^{-1}$, where $T_{\mathrm{1,ext}}(\omega)=1/\gamma_{\mathrm{q}}(\omega)$ is the contribution from just the external coupling to the drive line, $T_{\mathrm{1,diel}}(\omega)$ corresponds to the dielectric losses of the qubit, and $T_{\mathrm{1,oth}}(\omega)$ accounts for other sources of relaxation such as readout resonators, flux lines, etc.
@@
Figure 2c shows that within the filter stopband centered at 5 GHz, the transmission from the qubit port to the transmission line lies below $-110$ dB, greatly suppressed from $-72$ dB obtained for the standard drive line. Importantly, within the subharmonic frequency range of 1–2 GHz, we observe an increase of over 30 dB in the transmission of the on-chip filters over the standard case.
@@
In Fig. 2d, we examine the relaxation time $T_{\mathrm{1,ext}}$ of the qubit resulting from the external coupling to the drive line. Within the filter stopband, we observe an enhancement in $T_{\mathrm{1,ext}}$ by a factor of over 1000 with the $\lambda/4$ and $\lambda/2$ filters compared with the standard drive line. The estimated $T_{\mathrm{1,ext}}$ times at 5 GHz are as follows: 0.3 ms for the standard drive, 7161 ms for $\lambda/4$ filter, and  2421 ms for $\lambda/2$ filter. The widths of the frequency ranges where qubit $T_{\mathrm{1,ext}}$ exceeds 1 ms are 70 MHz and 450 MHz for $\lambda/4$ and $\lambda/2$ filter, respectively. We attribute this difference of feasible frequency ranges to the fact that the $\lambda/2$ filter is of higher order than the $\lambda/4$ filter since the coupling points of the qubit are located at the voltage anti-nodes where the gradient of the voltage with respect to position vanishes.
@@
Assuming a coupling that results in $T_{\mathrm{1,ext}}=1$ ms for a qubit driven resonantly through a standard transmission line, we estimate that a microwave pulse with a peak power of $-11$ dBm at room temperature is necessary to implement 10-ns single-qubit gates. With the above simulated parameters of the on-chip filters, corresponding peak powers of $-16$ dBm and $-22$ dBm will be required for $\lambda/4$ filter and $\lambda/2$ filter, respectively, using a subharmonic drive approach. Thus the filters combined with subharmonic driving may yield a significant reduction in power consumption for single-qubit gates. These values are calculated considering 60 dB of attenuation in the drive line and a simple rectangular pulse shape. Note that the peak powers may significantly increase for complex pulse shapes, which are currently being used to minimize the leakage out of the computational subspace for high-fidelity gates [27].
@@
Figure 3: Characterization of the on-chip filters. a,b Measured (dots) and simulated (dashed line) Rabi frequencies of the qubit for (a) $\lambda/4$ and (b) $\lambda/2$ filters as functions of the qubit frequency $f_{\mathrm{q}}=\omega_{\mathrm{q}}/(2\pi)$. We swept qubit frequency in the range of 4 to 6.5 GHz and carried out the Rabi measurement to characterize the coupling strength of the drive lines to the qubit. The dashed line represents the Rabi frequency estimated using $f_{\mathrm{R}}=\Omega_{\mathrm{R}}/2\pi=2\sqrt{\gamma_{\mathrm{q}}}\beta$, where $\gamma_{\mathrm{q}}=\mathrm{Re}[Y_{\mathrm{in}}]/C_{\mathrm{q}}$ and $\beta\propto\omega_{\mathrm{q}}^{-1/2}V$, where V is the peak-to-peak voltage amplitude at the chip level. Note that the measured Rabi frequency for the $\lambda/2$ filter shows two distinct stopbands centered around 4.7 GHz and 5.7 GHz, the presence of which is attributed to the unequal coupling at positions $x=0$ and $x=\mathrm{\lambda/2}$. c,d Measured (dots) and simulated (dashed line) relaxation time, $T_{1}$, of the qubit for (c) $\lambda/4$ and (d) $\lambda/2$ filters as functions of the qubit frequency. The dielectric loss tangent of $3\times 10^{-6}$ obtained from the fit is used to account for dielectric losses limiting the $T_{1}$, represented by the red dashed line. While orange dashed line corresponds to the case without the dielectric losses. The simulated $T_{1}$ also takes Purcell decay through the resonator into account. The error bars represent $1\sigma$ fitting uncertainty.
@@
We follow the typical steps to fabricate the device on a silicon wafer, which includes a niobium metallization layer on top. The micron-sized features in the niobium are created using maskless lithography and etching, while the Manhattan-style Josephson junctions are patterned using electron-beam lithography and thermal evaporation of aluminium. We cool down the device in a commercial Bluefors XLD500 dilution refrigerator for cryogenic measurements. For more detailed information on the fabrication process and the experimental setup, see the Supplementary Note S2 & S3 in Supplementary Information.
@@
We carry out typical time-domain measurements to characterize the device parameters summarized in Table S1 in Supplementary Information. To quantify the performance of the on-chip filters, we measured the Rabi frequency and the $T_{1}$ of the qubit around the target filter frequency of 5 GHz. A comprehensive description of the measurement sequence can be found in greater detail in the Supplementary Note S4 in Supplementary Information.
@@
Figures 3a, b show the measured Rabi frequency for qubits driven through the $\mathrm{\lambda/4}$ and $\mathrm{\lambda/2}$ filters. We observe an impressive 50-fold and 220-fold suppression in measured Rabi frequency at the stopbands of $\lambda/4$ and $\lambda/2$ filters, respectively, corresponding to 34 dB and 47 dB changes in $S_{21}$. The suppression factor is defined with respect to the maximum Rabi frequency observed in the qubit frequency range of 4–6.5 GHz. The simulated data depicted in Fig. 3a, b correspond to the Rabi frequency derived from the input admittance $Y_{\mathrm{in}}$ obtained through EM simulations, see Methods. Evidently, the measured and simulated Rabi frequencies exhibit a high level of agreement. Some minor deviations observed can be attributed to parasitic resonance and imperfections stemming from coupling with the control lines.
@@
Note the significant discrepancy in the measured Rabi frequency for the $\lambda/2$ filter shown in Fig. 3b in comparison to the simulated transmission depicted in Fig. 2c. In contrast to a single stopband at 5 GHz, we observe two stopbands centered around 4.7 GHz and 5.7 GHz. As detailed in Supplementary Note S1 in Supplementary Information, this splitting of the stopband mainly arises from the asymmetry in the coupling capacitance between the drive line and the qubit at the positions $x=0$ and $x=\lambda/2$, in addition to the fact that the coupling region at $x=\lambda/2$ is of finite length. Consequently, we show a good agreement between the measured and simulated results in Fig. 3b by introducing an asymmetric coupling at the open-end of the $\mathrm{\lambda/2}$ filter in the EM simulation.
@@
Figures 3c, d depict the measured $T_{1}$ for the $\mathrm{\lambda/4}$ and $\mathrm{\lambda/2}$ filter. In our device, we consistently observe a $T_{1}$ ceiling of approximately 10 $\mu\mathrm{s}$ across all qubits. To investigate the reason for this limited $T_{1}$, we fit the Purcell decay and dielectric-loss models to the measured $T_{1}$ of qubit Q1 near its readout resonator frequency, see Supplementary Note S5 in Supplementary Information. From the fit, we obtain a dielectric loss tangent of $3\times 10^{-6}$, confirming dielectric loss as the primary loss mechanism in our device. Consequently, we used the dielectric loss tangent of $3\times 10^{-6}$ in the EM simulations, yielding the simulation results for qubits Q2 and Q3 in good agreement with the experiments. Without considering dielectric losses in the simulation, the extrapolated $T_{1}$ at the filter frequency reaches hundreds of milliseconds range. Due to the Purcell decay through the resonator, the simulated $T_{1}$ at the second stopband centered around 5.7 GHz shows only a few milliseconds.
@@
For the qubit $T_{1}$ measured at the stopband of the filter, we observed a remarkable 20-fold improvement for the $\mathrm{\lambda/4}$ filter and an even more substantial 200-fold improvement for the $\mathrm{\lambda/2}$ filter in comparison to the values measured at the flux sweet-spot of the qubit, see Supplementary Table S1 in Supplementary Information. This enhancement has the potential for further improvements through the use of low-loss dielectric materials and implementation of state-of-the-art fabrication processes for high coherence.
@@
Figure 4: Spectroscopy and Rabi oscillations of qubit Q3 with resonant driving and subharmonic driving through the $\mathrm{\lambda/2}$ filter. a,b Normalized readout signal of the qubit as a function of the flux bias and the frequency of the (a) resonant drive and (b) subharmonic drive. The top diagonal feature indicates the lowest transition frequency of the qubit and the bottom feature reveals the two-photon transition to the second excited state of the qubit. c,d Normalized readout signal of the qubit as a function of the driving-voltage amplitude at room temperature and the frequency of the (c) resonant drive and (d) subharmonic drive for a flux bias of 0.32 $\Phi_{\mathrm{0}}$. Note the power broadening and up to $-70$ MHz alternating-current (AC) Stark shift in (d) with increasing pulse amplitude. e,f Normalized readout signal of the qubit as a function of the pulse length and the frequency of the (e) resonant drive and (f) subharmonic drive for a flux bias of 0.32 $\Phi_{\mathrm{0}}$. In (e), the pulse amplitude 1 V yields a Rabi frequency of 17.57 MHz and in (f), we have 0.9 V and 31.12 MHz, respectively, at the Stark-shifted frequency of 1.809 GHz.
Figures 4a,b present the flux spectroscopy results for qubit Q3 coupled to $\lambda/2$ filter using both resonant and subharmonic driving techniques. By sweeping the flux bias $\Phi$ threading through the SQUID in the units of flux quanta $\Phi_{\mathrm{0}}$, we tune the frequency of qubit Q3 in a range of 500 MHz near the second stopband of the $\lambda/2$ filter centered around 5.7 GHz. With resonant driving, we observe a major reduction in the qubit linewidth as it approaches the stopband at 5.7 GHz. Note that also the two-photon process to excite the qubit from its ground state to the second excited state is visible at different flux bias but frequency matching to 5.7 GHz, consistent with the filter protecting the qubit at its stopband rather than the qubit itself being substantially different at a certain flux bias. However, with subharmonic driving, we observe a broadened qubit linewidth at around $f=5.7/3=1.9$ GHz. This observation indicates that the $\lambda/2$ filter exhibits strong coupling at the subharmonic frequency but very weak coupling at the resonance frequency, which agrees with our intended design and purpose.
@@
Figures 4e,f illustrate Rabi oscillations of the qubit under resonant and subharmonic driving. For room-temperature pulse amplitudes of 1 V and 0.9 V, the measured Rabi frequencies for resonant and subharmonic driving are 17.57 MHz and 31.12 MHz, respectively. In the case of subharmonic drive, the Rabi frequency corresponding to the slowest oscillation occurs with an AC Stark shift of $-57.67$ MHz.